# Magnetically-Driven Neutron-Rich Ejecta Unleashed: Global 3D Neutrino-GRMHD Simulations of Collapsars Reveal the Conditions for $r$-process Nucleosynthesis

### v1

@@ 3.2 Resolving the magnetorotational instability (MRI)
Even though our initial magnetic field is not dynamically important, upon the disk formation the magnetorotational instability is well-resolved (MRI; Balbus & Hawley, 1991) in all models (except for pB0). In contrast, when MRI is not sufficiently resolved, it can lead to a runaway density buildup at the inner boundary of the disk, leading to spuriously low values of $\,{Y_{\rm e}}$. Our hydrodynamic model (pB0) is a prime example of this issue on a global scale (see also Dean & Fernández, 2024a, for a similar trade-off in hydrodynamic flows); locally, this can occur if the unresolved region is large enough.
@@ 3.3 Equation of state
For the equation of state (EOS), we adopt Helmholtz EOS, following the approach of Timmes & Swesty (2000), where thermodynamic quantities such as pressure, specific internal energy, and entropy, are computed as the sum of contributions from an ideal gas of ions, radiation, and electrons/positrons of arbitrary degeneracy:
@@
The first two components on the right hand side are calculated analytically. For the degeneracy pressure, we use a pre-computed tabulated form of the Helmholtz free energy and its derivatives as functions of density and temperature. The table is spaced logarithmically in density (541 points, $\rho\in[10^{-12},10^{15}]\,\rm{g\,cm}^{-3}$), and temperature (201 points, $T\in[10^{3},10^{13}]\,\rm{K}$). It is linear in the electron fraction ($\,{Y_{\rm e}}\in[0,1]$). We note that at densities approaching nuclear saturation density ($\sim 10^{14}\rm{g/cm^{3}}$), this EOS is no longer valid, since degeneracy pressure is not the main source of pressure in this regime. In practice, in our simulations densities do not exceed $\approx 10^{13}\rm{g/cm^{3}}$, well within the validity range of the Helmholtz EOS.
3.4 Neutrino transport implementation
Neutrinos play a crucial role in setting the composition of the accretion disk and its outflows. The degree of neutron-richness, expressed through electron fraction $\,{Y_{\rm e}}$, will determine whether $r$-process occurs in the outflows, alongside the entropy and the expansion timescale (e.g., Lippuner & Roberts, 2015). Therefore, we need accurate neutrino transport to model the compositional evolution of the ejected material. In this work, we implement a two-moment (M1) scheme for the neutrino transport, following Foucart et al. (2015, 2016); Sądowski et al. (2014) and McKinney et al. (2014), where neutrino-matter interactions are included via a lookup table (Nulib, O’Connor, 2015). Here, we present a brief overview of the scheme and its implementation. We provide a more detailed description of the implementation details, numerical tests, and calculations of the neutrino and antineutrino properties in Appendix A.
We evolve 3 neutrino species $(\nu_{\rm e},\bar{\nu}_{\rm e},\nu_{\rm x})$, where the latter is the combination of 4 species $(\nu_{\mu},\bar{\nu}_{\mu},\nu_{\tau},\bar{\nu}_{\tau})$. For a single species, we write the neutrino energy-momentum tensor as:
@@ 3.4 Neutrino transport implementation
where $u^{\mu}$ is the 4-velocity and $g^{\mu\nu}$ is the contravariant metric, $(E,F^{\mu},P^{\mu\nu})$, $(J,H^{\mu},L^{\mu\nu})$ are the moments of the neutrino distribution function (neutrino energy, fluxes and stress tensor) measured by an observer in the lab (coordinate) and the fluid frames respectively. We introduce a “radiation” frame that moves at 4-velocity, $u^{\mu}_{\rm R}$, whose observer measures the energy density, $E_{\rm R}$. The evolution equations come from the energy-momentum conservation equation:
@@
where $\rho$ is the gas density, $u_{\rm g},P_{\rm g}$ are the gas internal energy and pressure, all measured in the rest frame of the gas. Here, $b^{\mu}$ is the contravariant magnetic field 4-vector (and $b^{2}=b^{\mu}b_{\mu}$).
To close the evolution equations we assume a closure relation expressing the stress tensor in terms of the energy and flux densities. In our work, we follow an approach by Sądowski et al. (2014); McKinney et al. (2014) where we adopt a closure by Levermore (1984), assuming that there exists a “radiation” frame, in which neutrino radiation is isotropic. Then, the lab frame moments can be found by boosting the radiation stress-energy tensor from the radiation frame to the lab (coordinate) frame (with velocity $-u^{\mu}_{\rm R}$).
@@ 4 Simulation results
Figure 1: Models with flat density cores (cB) develop large values of mass accretion rate, $\dot{M}\sim 10\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$ [panel (a)], and dimensionless magnetic flux [panel (b)] that are favorable for the production of neutron-rich material and outflows (see Fig. 2). Once the dimensionless magnetic flux exceeds the critical value, $\phi\sim 10$ [panel (b)], the jet launching efficiency [panel (c)] and power [panel (d)] surge. Spurred by high $\dot{M}$, the ignition radius, $R_{\rm ign}$ [panel (e)], initially rises, but drops as $\phi$ increases [panel (b)] and obstructs $\dot{M}$ [panel (a)]. Circles and squares mark $t_{\rm jet}$ and $t_{\rm MAD}$ of each model, respectively. To reduce clutter, we smoothed all quantities over $\Delta t=10^{3}\,{r_{\rm g}}/c=0.02$ s.
Figure 1(a) illustrates the time evolution of mass accretion rate measured at $r=5\,{r_{\rm g}}$ (to avoid contamination by density floors at smaller distances) across all models. The mass density profile of the innermost shells dictates the mass accretion rate. We consider flat core density profiles with $\alpha_{\rm p}=0$ in models cBs, cBw, cBwA0, as expected at the initial time of collapse in GRB progenitors (e.g., Woosley & Bloom, 2006). However, in our simulations, $t=0$ marks the time of BH formation, which is preceded by a few seconds-long proto-NS phase (e.g., Aloy & Obergaulinger, 2021). During this phase, the innermost shells exhibit a free-fall radial density profile, $\alpha_{\rm p}=1.5$, which we adopt in the models pBs, pBw, pB0  (Halevi et al., 2023). To maintain the same total mass in the star, these models feature higher density values at the center than those with $\alpha_{\rm p}=0$.
Higher central densities lead to higher mass accretion rates that evolve as $\dot{M}\sim t^{1-2\alpha_{\rm p}/3}$ (see Gottlieb et al., 2022a). Thus, in models with a flat core (models cBw, cBs, cBwA0), the mass accretion rate is initially very high ($\dot{M}\sim 10\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$) but begins to decline already at $t\lesssim 0.1\,\,{{\rm s}}$. Conversely, models with $\alpha_{\rm p}=1.5$ (models pBw, pB0) exhibit a lower but steadier mass accretion rate, which at $t\gtrsim 1\,\,{{\rm s}}$ surpasses that of the flat core models. When strong magnetized outflows form (model pBs; green line), feedback from these outflows hinders accretion, leading to a decline in the mass accretion rate regardless of the density profile.
Figure 1(b,c) shows the dimensionless magnetic flux on the event horizon (Eq. 8) and jet launching efficiency:
@@
where the jet power (or outflow power for model cBwA0) is defined as the power contained in magnetically-dominated ($\sigma=b^{2}/\rho>1$) outflows:
@@
Here, $T^{r}_{t}$ gives the negative of the radial energy flux density, a component of the mixed stress-energy tensor, $T^{\mu}_{\nu}$. The last term in parentheses removes the contribution of the rest-mass energy flux from the total energy flux. When the dimensionless magnetic flux exceeds $\phi\approx 10$ [panel (b)], a relativistic jet launches, as indicated by a sharp rise in the jet efficiency, $\eta_{\rm jet}$ [panel (c)]. This effect is most pronounced in models pBs, cBs, cBw, where once $\phi\gtrsim 10$ at $t\lesssim 0.1\,\,{{\rm s}}$, the jet efficiency surges from $\eta_{\rm jet}<10^{-3}$ to $\eta_{\rm jet}\gtrsim 0.1$. This result is consistent with an order of magnitude estimate of the critical $\phi$ for jet launching in §2. Ultimately, the dimensionless magnetic flux reaches the asymptotic MAD level of $\phi\approx 50$ [marked by a grey dashed line in Fig. 1(b)], where the jet launching efficiency stabilizes at an order unity. We note that the critical $\phi$ for magnetized outflow launching depends on the BH spin, as seen in model cBwA0. In this model, the outflows only launch after the MAD state is established, with the launching efficiency remaining low at $\eta_{\rm jet}\lesssim 10^{-2}$.
@@
Figure 2: Density-weighted average electron fraction as a function of radius at various times for different models. The regions with neutron-rich material closely track the ignition radii (white lines). The increase in disk magnetic flux leads to jet launching (yellow dashed lines) and a transition to MAD (red dashed lines). This transition signifies the drop in the accretion timescale, which ultimately ends the neutron-rich phase of the disk.
@@
The product of the accretion power and the jet launching efficiency determines the jet power, depicted in Fig. 1(d). In most models, the combination of high efficiency and high mass accretion rates initially produces extremely powerful jets. However, once $\phi\approx 50$, the jet power diminishes over time, following the decline in the mass accretion rate. In model cBwA0, the non-spinning BH system can only launch outflows by extracting the angular momentum from the disk (not the BH). Consequently, the outflow power in this model is lower than that in rest of the magnetized models.
@@
Figure 3: Meridional cuts of the mass density (left-hand side) and electron fraction (right-hand side) for all models (different columns). The top row displays the system before jet launching ($t<t_{\rm jet}$), featuring neutron-rich disks where $\,{Y_{\rm e}}$ varies with mass accretion rate and disk magnetization. After the launching of jets and disk winds ($t_{\rm jet}<t<t_{\rm MAD}$), neutron-rich material is ejected from the disk midplane and carried by the shocked cocoon that engulfs the jets. In the bottom row, magnetic fields are dynamically important ($t>t_{\rm MAD}$), shortening the accretion timescale, which leads to disk deneutronization and the cessation of neutron-rich outflows.
@@
Fig. 1(e) displays the time evolution of the radius within which neutrino cooling is efficient, $R_{\rm ign}$. This radius increases as the cooling timescale grows and the accretion timescale shortens. In our magnetized simulations, the accretion timescale remains small due to efficient angular momentum transport, driven by large-scale magnetic fields. For models with mass accretion rates below $\dot{M}\sim 1\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$ (pBs, pBw), cooling is inefficient due to lower density, resulting in $R_{\rm ign}\sim 10\,{r_{\rm g}}$. Conversely, in models with very high mass accretion rates (cBw, cBw, cBwA0), increased density leads to reduced cooling timescales and a larger $R_{\rm ign}\sim 100\,{r_{\rm g}}$. In all cases, the accretion timescale plummets once the MAD state is established. Combined with the decrease in mass accretion rates, which signify longer cooling timescales, $R_{\rm ign}$ inevitably drops. Hence, the MAD state marks the transition from the neutron-rich disk phase to the neutron-poor disk phase, accompanied by an increase in $\,{Y_{\rm e}}$.
@@
Figure 2 demonstrates the role of $R_{\rm ign}$ in estimating the electron fraction in the disk, delineating the density-weighted angle-averaged electron fraction $\langle\,{Y_{\rm e}}\rangle_{\rho}^{\theta,\varphi}(r,t)$ (Eq. 7) as a function of radius and time. The definition of $R_{\rm ign}$ in Eq. (4) is supported by a good agreement between $R_{\rm ign}$ (solid white line in Fig. 2) and the average radial extents of the neutron-rich material in the disk across all models. By the time of jet (outflow) launching (yellow vertical dashed lines in models pBs, cBs, cBw, cBwA0), some disk material is expelled via disk winds. This ejection contributes to a rise in the average $\,{Y_{\rm e}}$. The onset of the MAD state (red vertical dashed lines) marks the end of the neutron-rich disk state. If the magnetic flux on the BH is insufficient for outflow launching, the neutron-rich material is swiftly accreted (as in model pBw) or a long-lived neutron-rich disk is established (model pB0).
@@
Figure 4: 3D rendering of the $\,{Y_{\rm e}}$ in the ejecta of model cBw, at different times. Around the time of jet launching (left panel), the disk is neutron-rich (low $\,{Y_{\rm e}}$). At $t=0.2$ s, the jet and neutron-rich winds are launched, and their interaction forms a neutron-rich cocoon. As the cocoon interacts with the neutron-poor gas in the star, significant mixing occurs, increasing the cocoon electron fraction (middle panels). At later times (right panel), the disk density drops significantly, resulting in higher $\,{Y_{\rm e}}$, while the outflows continue to mix with the star material, further raising $\,{Y_{\rm e}}$ in the cocoon.
@@
Figure 3 shows meridional 2D slices of mass density (left-hand-side of each panel) and electron fraction $\,{Y_{\rm e}}$ (right-hand-side). The first row shows the disk evolution prior to jet launching, at $t<t_{\rm jet}$. Jet (outflow) launch time $t_{\rm jet}$ is used to mark the moment when outflow launching efficiency exceeds $\eta_{\rm jet}\gtrsim 10^{-3}$ (or when $\phi\lesssim 10$) and is used for illustrative purposes (as opposed to $t_{\rm MAD}$ defined by the moment when $\phi\approx 50$). In this phase, $\dot{M}>\dot{M}_{\rm ign}$, which results in the electron fraction dropping below $\,{Y_{\rm e}}=0.5$. The specific $\,{Y_{\rm e}}$ value is governed by the ratio of $t_{\rm cool}/t_{\rm acc}$, which is, in turn, dictated by the mass accretion rate and magnetic flux. In models with very high accretion rates $\dot{M}\sim 10\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$ (cBs, cBw, cBwA0), where $R_{\rm ign}$ is large, the electron fraction in the disk drops to $\,{Y_{\rm e}}\lesssim 0.2$. In models with moderate accretion rates and appreciable magnetic fields (pBw, pBs), $R_{\rm ign}$ is smaller and $\,{Y_{\rm e}}\approx 0.3$. For comparison, in the hydrodynamic model (pB0; rightmost panels), the absence of MRI-driven turbulent torques results in a longer accretion timescale. Consequently, the gas has enough time to cool efficiently, allowing the electron fraction to drop as low as $\,{Y_{\rm e}}\approx 0.1$, even with a moderate mass accretion rate. Without powerful outflows and efficient accretion, the disk remains neutron-rich until the end of the simulation, while slowly spreading as its mass increases, as indicated by the rising $R_{\rm ign}$ [Fig. 1(e)].
@@
The second row of Fig. 3 depicts the system shortly after jet launching (except for models pBw, pB0  where no jets are present), at $t_{\rm jet}<t<t_{\rm MAD}$ ($10\lesssim\phi\lesssim 50$). The growth of $\phi$ in the disk drives both the jet launching from the BH and emergence of disk winds. The disk winds can eject portions of the low $\,{Y_{\rm e}}$ material from the disk midplane. As the jets ($\,{Y_{\rm e}}\approx 0.5$) shock the neutron-rich winds, they inflate a pressurized neutron-rich cocoon, as shown by the low $\,{Y_{\rm e}}$ values observed in the jet wings. For $a=0$ (model cBwA0), the absence of relativistic jets implies that the magnetized disk winds are the leading force in carrying the neutron-rich ejecta to the stellar edge.
@@
The bottom row of Fig. 3, which shows the systems at $t>t_{\rm MAD}$, depicts the transition from a neutron-rich to a neutron-poor disk once a MAD state is reached. Namely, the increased magnetic flux and reduced mass accretion rate lower $R_{\rm ign}$, halting the production of neutron-rich material in the disk. Concurrently, the electron fraction in the cocoon rises over time due to two factors. First, the deneutronization of the disk prevents the cocoon from replenishing its neutron-rich gas. Second, the interaction between the cocoon’s bulk and the infalling neutron-poor gas increases its $\,{Y_{\rm e}}$. Consequently, while the jet launching marks the efficient ejection of neutron-rich material from the disk, it also signifies the onset of increasing $\,{Y_{\rm e}}$. In model pBw, where the amount of magnetic flux is insufficient to launch outflows, neutron-rich material remains in the disk and eventually accretes onto the BH.
@@
As low $\,{Y_{\rm e}}$ matter is ejected, weak interactions, particularly $e^{\pm}$ pair captures, freeze out and the composition is then governed by either neutrino absorption or mixing. Unlike in compact binary mergers, the ejecta continues to interact with the dense star, allowing for ongoing compositional changes through mixing. Figure 5 depicts unbound ejecta histograms in $\,{Y_{\rm e}}$ at various times across all models with ejecta. The unbound mass is defined through the Bernoulli criterion:
@@
where $u_{t}$ is the covariant time component of the four-velocity and $u_{g}$ is the thermal energy density. Before jet launching, $b^{2}/\rho$ rises above unity along the polar regions where the gas free-falls. However, this magnetized gas ends up being accreted onto the BH. Therefore, we disregard the magnetic field contribution for the unbound criterion at $t<T_{\rm jet}$.

### v2

@@ 3.2 Resolving the magnetorotational instability (MRI)
Even though our initial magnetic field is not dynamically important, upon the disk formation, the magnetorotational instability is well resolved (MRI; Balbus & Hawley, 1991) in all models (except for pB0). In contrast, when MRI is not sufficiently resolved, it can lead to a runaway density buildup at the inner boundary of the disk, leading to spuriously low values of $\,{Y_{\rm e}}$. Our hydrodynamic model (pB0) is a prime example of this issue on a global scale (see also Dean & Fernández, 2024a, for a similar trade-off in hydrodynamic flows); locally, this can occur if the unresolved region is large enough.
@@ 3.3 Equation of state
For the EOS, we adopt Helmholtz EOS, following the approach of Timmes & Swesty (2000), where thermodynamic quantities such as pressure, specific internal energy, and entropy are computed as the sum of contributions from an ideal gas of ions, radiation, and electrons/positrons of arbitrary degeneracy:
@@
The first two components on the right-hand side are calculated analytically. For the degeneracy pressure, we use a precomputed tabulated form of the Helmholtz free energy and its derivatives as functions of density and temperature. The table is spaced logarithmically in density (541 points, $\rho\in[10^{-12},10^{15}]\,\rm{g\,cm}^{-3}$) and temperature (201 points, $T\in[10^{3},10^{13}]\,\rm{K}$). It is linear in the electron fraction ($\,{Y_{\rm e}}\in[0,1]$). We note that at densities approaching nuclear saturation density ($\sim 10^{14}\rm{g/cm^{3}}$), this EOS is no longer valid, since degeneracy pressure is not the main source of pressure in this regime. In practice, in our simulations, densities do not exceed $\approx 10^{13}\rm{g/cm^{3}}$, well within the validity range of the Helmholtz EOS.
3.4 Neutrino-transport implementation
Neutrinos play a crucial role in setting the composition of the accretion disk and its outflows. The degree of neutron richness, expressed through electron fraction $\,{Y_{\rm e}}$, will determine whether $r$-process occurs in the outflows, alongside the entropy and the expansion timescale (e.g., Lippuner & Roberts, 2015). Therefore, we need accurate neutrino transport to model the compositional evolution of the ejected material. In this work, we implement an M1 scheme for the neutrino transport, following Foucart et al. (2015, 2016); Sądowski et al. (2014) and McKinney et al. (2014), where neutrino-matter interactions are included via a lookup table (Nulib, O’Connor, 2015). Here, we present a brief overview of the scheme and its implementation. We provide a more detailed description of the implementation details, numerical tests, and calculations of the neutrino and antineutrino properties in Appendix A.
We evolve three neutrino species $(\nu_{\rm e},\bar{\nu}_{\rm e},\nu_{\rm x})$, where the latter is the combination of four species $(\nu_{\mu},\bar{\nu}_{\mu},\nu_{\tau},\bar{\nu}_{\tau})$. For a single species, we write the neutrino energy-momentum tensor as
@@ 3.4 Neutrino-transport implementation
where $u^{\mu}$ is the 4-velocity and $g^{\mu\nu}$ is the contravariant metric; and $(E,F^{\mu},P^{\mu\nu})$ and $(J,H^{\mu},L^{\mu\nu})$ are the moments of the neutrino distribution function (neutrino energy, fluxes, and stress tensor) measured by an observer in the lab (coordinate) and the fluid frames, respectively. We introduce a “radiation” frame that moves at 4-velocity, $u^{\mu}_{\rm R}$, whose observer measures the energy density, $E_{\rm R}$. The evolution equations come from the energy-momentum conservation equation:
@@
where $\rho$ is the gas density and $u_{\rm g},P_{\rm g}$ are the gas internal energy and pressure, all measured in the rest frame of the gas. Here, $b^{\mu}$ is the contravariant magnetic field 4-vector (and $b^{2}=b^{\mu}b_{\mu}$).
To close the evolution equations, we assume a closure relation expressing the stress tensor in terms of the energy and flux densities. In our work, we follow an approach by Sądowski et al. (2014) and McKinney et al. (2014), where we adopt a closure by Levermore (1984), assuming that there exists a “radiation” frame, in which neutrino radiation is isotropic. Then, the lab frame moments can be found by boosting the radiation stress-energy tensor from the radiation frame to the lab (coordinate) frame (with velocity $-u^{\mu}_{\rm R}$).
@@ 4 Simulation results
Figure 1: Models with flat density cores (cB) develop large values of mass accretion rate, $\dot{M}\sim 10\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$ (panel (a)), and dimensionless magnetic flux [panel (b)] that are favorable for the production of neutron-rich material and outflows (see Fig. 2). Once the dimensionless magnetic flux exceeds the critical value, $\phi\sim 10$ (panel (b)), the jet-launching efficiency (panel (c)) and power (panel (d)) surge. Spurred by high $\dot{M}$, the ignition radius, $R_{\rm ign}$ (panel (e)), initially rises but drops as $\phi$ increases (panel (b)) and obstructs $\dot{M}$ (panel (a)). Circles and squares mark $t_{\rm jet}$ and $t_{\rm MAD}$ of each model, respectively. To reduce clutter, we smoothed all quantities over $\Delta t=10^{3}\,{r_{\rm g}}/c=0.02$ s.
Figure 1(a) illustrates the time evolution of mass accretion rate measured at $r=5\,{r_{\rm g}}$ (to avoid contamination by density floors at smaller distances) across all models. The mass density profile of the innermost shells dictates the mass accretion rate. We consider flat core density profiles with $\alpha_{\rm p}=0$ in models cBs, cBw, and cBwA0, as expected at the initial time of collapse in GRB progenitors (e.g., Woosley & Bloom, 2006). However, in our simulations, $t=0$ marks the time of BH formation, which is preceded by a few seconds long proto-NS phase (e.g., Aloy & Obergaulinger, 2021). During this phase, the innermost shells exhibit a free-fall radial density profile, $\alpha_{\rm p}=1.5$, which we adopt in the models pBs, pBw, and pB0  (Halevi et al., 2023). To maintain the same total mass in the star, these models feature higher density values at the center than those with $\alpha_{\rm p}=0$.
Higher central densities lead to higher mass accretion rates that evolve as $\dot{M}\sim t^{1-2\alpha_{\rm p}/3}$ (see Gottlieb et al., 2022a). Thus, in models with a flat core (models cBw, cBs, and cBwA0), the mass accretion rate is initially very high ($\dot{M}\sim 10\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$) but begins to decline already at $t\lesssim 0.1\,\,{{\rm s}}$. Conversely, models with $\alpha_{\rm p}=1.5$ (models pBwand pB0) exhibit a lower but steadier mass accretion rate, which at $t\gtrsim 1\,\,{{\rm s}}$ surpasses that of the flat core models. When strong magnetized outflows form (model pBs; green line), feedback from these outflows hinders accretion, leading to a decline in the mass accretion rate regardless of the density profile. We note that in the model without magnetic fields, pB0, most of the accretion occurs through the polar funnel and not through the disk, leading to elevated $\dot{M}$ in comparison with models pBsand pBw.
Figure 1(b), (c) show the dimensionless magnetic flux on the event horizon (Eq. 8) and jet-launching efficiency:
@@
where the jet power (or outflow power for model cBwA0) is defined as the power contained in magnetically dominated ($\sigma=b^{2}/\rho>1$) outflows:
@@
Here, $T^{r}_{t}$ gives the negative of the radial energy flux density, a component of the mixed stress-energy tensor, $T^{\mu}_{\nu}$. The last term in parentheses removes the contribution of the rest-mass energy flux from the total energy flux. When the dimensionless magnetic flux exceeds $\phi\approx 10$ (panel (b)), a relativistic jet launches, as indicated by a sharp rise in the jet efficiency, $\eta_{\rm jet}$ (panel (c)). This effect is most pronounced in models pBs, cBs, and cBw, where once $\phi\gtrsim 10$ at $t\lesssim 0.1\,\,{{\rm s}}$, the jet efficiency surges from $\eta_{\rm jet}<10^{-3}$ to $\eta_{\rm jet}\gtrsim 0.1$. This result is consistent with an order-of-magnitude estimate of the critical $\phi$ for jet launching in §2. Ultimately, the dimensionless magnetic flux reaches the asymptotic MAD level of $\phi\approx 50$ (marked by a gray dashed line in Fig. 1(b)), where the jet-launching efficiency stabilizes at an order unity. We note that the critical $\phi$ for magnetized outflow launching depends on the BH spin, as seen in model cBwA0. In this model, the outflows only launch after the MAD state is established, with the launching efficiency remaining low at $\eta_{\rm jet}\lesssim 10^{-2}$.
@@
Figure 2: Density-weighted average electron fraction as a function of radius at various times for different models. The regions with neutron-rich material closely track the ignition radii (white lines). The increase in disk magnetic flux leads to jet launching (yellow dashed lines) and a transition to MAD (red dashed lines). This transition signifies the drop in the accretion timescale and increase in the neutrino-cooling timescale due to decreasing densities, which ultimately ends the neutron-rich phase of the disk.
@@
The product of the accretion power and the jet-launching efficiency determines the jet power, depicted in Fig. 1(d). In most models, the combination of high efficiency and high mass accretion rates initially produces extremely powerful jets. However, once $\phi\approx 50$, the jet power diminishes over time, following the decline in the mass accretion rate. In model cBwA0, the nonspinning BH system can only launch outflows by extracting the angular momentum from the disk (not the BH). Consequently, the outflow power in this model is lower than that in rest of the magnetized models.
@@
Figure 3: Meridional cuts of the mass density (left-hand side) and electron fraction (right-hand side) for all models (different columns). The top row displays the system before jet launching ($t<t_{\rm jet}$), featuring neutron-rich disks where $\,{Y_{\rm e}}$ varies with mass accretion rate and disk magnetization. After the launching of jets and disk winds ($t_{\rm jet}<t<t_{\rm MAD}$), neutron-rich material is ejected from the disk midplane and carried by the shocked cocoon that engulfs the jets. In the bottom row, magnetic fields are dynamically important ($t>t_{\rm MAD}$), shortening the accretion timescale and reducing the disk densities, leading to a growing neutrino-cooling timescale, which leads to disk deneutronization and the cessation of neutron-rich outflows.
@@
Fig. 1(e) displays the time evolution of the radius within which neutrino cooling is efficient, $R_{\rm ign}$. This radius increases as the cooling timescale grows and the accretion timescale shortens. In our magnetized simulations, the accretion timescale remains small due to efficient angular momentum transport, driven by large-scale magnetic fields. For models with mass accretion rates below $\dot{M}\sim 1\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$ (pBsand pBw), cooling is inefficient due to lower density, resulting in $R_{\rm ign}\sim 10\,{r_{\rm g}}$. Conversely, in models with very high mass accretion rates (cBw, cBw, and cBwA0), increased density leads to reduced cooling timescales and a larger $R_{\rm ign}\sim 100\,{r_{\rm g}}$. In all cases, the accretion timescale plummets once the MAD state is established. Combined with the decrease in mass accretion rates, which reflect lower disk densities, which signify longer cooling timescales (see Fig. 13), $R_{\rm ign}$ inevitably drops. Hence, the MAD state marks the transition from the neutron-rich disk phase to the neutron-poor disk phase, accompanied by an increase in $\,{Y_{\rm e}}$.
@@
Figure 2 demonstrates the role of $R_{\rm ign}$ in estimating the electron fraction in the disk, delineating the density-weighted angle-averaged electron fraction $\langle\,{Y_{\rm e}}\rangle_{\rho}^{\theta,\varphi}(r,t)$ (Eq. 7) as a function of radius and time. The definition of $R_{\rm ign}$ in Eq. (4) is supported by a good agreement between $R_{\rm ign}$ (solid white line in Fig. 2) and the average radial extents of the neutron-rich material in the disk across all models. By the time of jet (outflow) launching (yellow vertical dashed lines in models pBs, cBs, cBw, and cBwA0), some disk material is expelled via disk winds. This ejection contributes to a rise in the average $\,{Y_{\rm e}}$. The onset of the MAD state (red vertical dashed lines) marks the end of the neutron-rich disk state. If the magnetic flux on the BH is insufficient for outflow launching, the neutron-rich material is swiftly accreted (as in model pBw), or a long-lived neutron-rich disk is established (model pB0).
@@
Figure 4: 3D rendering of the $\,{Y_{\rm e}}$ in the ejecta of model cBw, at different times. Around the time of jet launching (left panel), the disk is neutron rich (low $\,{Y_{\rm e}}$). At $t=0.2$ s, the jet and neutron-rich winds are launched, and their interaction forms a neutron-rich cocoon. As the cocoon interacts with the neutron-poor gas in the star, significant mixing occurs, increasing the cocoon electron fraction (middle panels). At later times (right panel), the disk density drops significantly, resulting in higher $\,{Y_{\rm e}}$, while the outflows continue to mix with the star material, further raising $\,{Y_{\rm e}}$ in the cocoon. An animation of the evolution of $\,{Y_{\rm e}}$ in the ejecta of model cBw  is available here. The animated 3D rendering is similar to the static version, but it uses a different color table and jet orientation. The animated version of the cBw  model proceeds from $t=0.0$ to $1.02$ s over the course of the 54 s movie.
@@
Figure 3 shows meridional 2D slices of mass density (left-hand side of each panel) and electron fraction $\,{Y_{\rm e}}$ (right-hand side). The first row shows the disk evolution prior to jet launching, at $t<t_{\rm jet}$. Jet (outflow) launch time $t_{\rm jet}$ is used to mark the moment when outflow launching efficiency exceeds $\eta_{\rm jet}\gtrsim 10^{-3}$ (or when $\phi\lesssim 10$) and is used for illustrative purposes (as opposed to $t_{\rm MAD}$ defined by the moment when $\phi\approx 50$). In this phase, $\dot{M}>\dot{M}_{\rm ign}$, which results in the electron fraction dropping below $\,{Y_{\rm e}}=0.5$. The specific $\,{Y_{\rm e}}$ value is governed by the ratio of $t_{\rm cool}/t_{\rm acc}$, which is, in turn, dictated by the mass accretion rate and magnetic flux. In models with very high accretion rates $\dot{M}\sim 10\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$ (cBs, cBw, and cBwA0), where $R_{\rm ign}$ is large, the electron fraction in the disk drops to $\,{Y_{\rm e}}\lesssim 0.2$. In models with moderate accretion rates and appreciable magnetic fields (pBwand pBs), $R_{\rm ign}$ is smaller and $\,{Y_{\rm e}}\approx 0.3$. For comparison, in the hydrodynamic model (pB0; rightmost panels), the absence of MRI-driven turbulent torques results in a longer accretion timescale. Consequently, the gas has enough time to cool efficiently, allowing the electron fraction to drop as low as $\,{Y_{\rm e}}\approx 0.1$, even with a moderate mass accretion rate. Without powerful outflows and efficient accretion, the disk remains neutron rich until the end of the simulation while slowly spreading as its mass increases, as indicated by the rising $R_{\rm ign}$ (Fig. 1(e)).
@@
The second row of Fig. 3 depicts the system shortly after jet launching (except for models pBwand pB0  where no jets are present), at $t_{\rm jet}<t<t_{\rm MAD}$ ($10\lesssim\phi\lesssim 50$). The growth of $\phi$ in the disk drives both the jet launching from the BH and emergence of disk winds. The disk winds can eject portions of the low-$\,{Y_{\rm e}}$ material from the disk midplane. As the jets ($\,{Y_{\rm e}}\approx 0.5$) shock the neutron-rich winds, they inflate a pressurized neutron-rich cocoon, as shown by the low $\,{Y_{\rm e}}$ values observed in the jet wings. For $a=0$ (model cBwA0), the absence of relativistic jets implies that the magnetized disk winds are the leading force in carrying the neutron-rich ejecta to the stellar edge.
@@
The bottom row of Fig. 3, which shows the systems at $t>t_{\rm MAD}$, depicts the transition from a neutron-rich to a neutron-poor disk once a MAD state is reached. Namely, the increased magnetic flux and reduced mass accretion rate lower $R_{\rm ign}$, halting the production of neutron-rich material in the disk. Concurrently, the electron fraction in the cocoon rises over time due to two factors. First, the deneutronization of the disk prevents the cocoon from replenishing its neutron-rich gas. Second, the interaction between the cocoon’s bulk and the infalling neutron-poor gas increases its $\,{Y_{\rm e}}$. Consequently, while the jet launching marks the efficient ejection of neutron-rich material from the disk, it also signifies the onset of increasing $\,{Y_{\rm e}}$. In model pBw, where the amount of the magnetic flux is insufficient to launch outflows, neutron-rich material remains in the disk and eventually accretes onto the BH.
@@
As low-$\,{Y_{\rm e}}$ matter is ejected, weak interactions, particularly $e^{\pm}$ pair captures, freeze-out and the composition is then governed by either neutrino absorption or mixing. Unlike in compact binary mergers, the ejecta continues to interact with the dense star, allowing for ongoing compositional changes through mixing. Figure 5 depicts unbound ejecta histograms in $\,{Y_{\rm e}}$ at various times across all models with ejecta. The unbound mass is defined through the Bernoulli criterion:
@@
where $u_{t}$ is the covariant time component of the 4-velocity and $u_{g}$ is the thermal energy density. Before jet launching, $b^{2}/\rho$ rises above unity along the polar regions where the gas free-falls. However, this magnetized gas ends up being accreted onto the BH. Therefore, we disregard the magnetic field contribution for the unbound criterion at $t<T_{\rm jet}$.